OpenStudy (anonymous):

I WILL GIVE MEDAL FOR BEST EXPLANATION!!! How to explain Negative Numbers with exponents?

OpenStudy (anonymous):

I need if with the number negative not the exponent but thanks anyway it was still helpful

OpenStudy (anonymous):

@awesomeness812

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

i mean yea

OpenStudy (anonymous):

I need if with the number negative not the exponent but thanks anyway it was still helpful @awesomeness812

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

... could you help with a negative number... not a negative exponent?

OpenStudy (anonymous):

(-x)^k if k=2n (pair) (-x)^k=x^k and if k=2n+1 (impair) (-x)^k=-(x^k) .